Stocks A and B have the following probability distributions of expected future returns:

Probability     A     B
0.1 (5 %) (37 %)
0.1 3 0
0.6 14 21
0.1 20 29
0.1 31 45
  1. Calculate the expected rate of return, , for Stock B ( = 13.30%.) Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. Calculate the standard deviation of expected returns, σA, for Stock A (σB = 20.55%.) Do not round intermediate calculations. Round your answer to two decimal places.

      %

    Now calculate the coefficient of variation for Stock B. Do not round intermediate calculations. Round your answer to two decimal places.

  3. Assume the risk-free rate is 3.5%. What are the Sharpe ratios for Stocks A and B? Do not round intermediate calculations. Round your answers to four decimal places.

    Stock A:

    Stock B:
Answer #1
a & b] STOCK A:
Probability [p] Rate of return [%]   [r] E[r] = p*r d = r-E[r] d^2 p*d^2
0.10 -5 -0.50 -18.30 334.89 33.49
0.10 3 0.30 -10.30 106.09 10.61
0.60 14 8.40 0.70 0.49 0.29
0.10 20 2.00 6.70 44.89 4.49
0.10 31 3.10 17.70 313.29 31.33
13.30 80.21
Expected return 13.30
Variance 80.21
SD = 80.21^0.5 = 8.96
Coefficient of variation = 8.96/13.3 = 0.67
STOCK B:
Probability [p] Rate of return [%]   [r] E[r] = p*r d = r-E[r] d^2 p*d^2
0.10 -37 -3.70 -53.30 2840.89 284.09
0.10 0 0.00 -16.30 265.69 26.57
0.60 21 12.60 4.70 22.09 13.25
0.10 29 2.90 12.70 161.29 16.13
0.10 45 4.50 28.70 823.69 82.37
16.30 422.41
Expected return 16.30
Variance 422.41
SD = 422.41^0.5 = 20.55
Coefficient of variation = 20.55/16.3 = 1.26
c] Sharpe ratio = [Expected return-Risk free rate]/Standard deviation
Stock A = (13.3-3.5)/8.96 = 1.09
Stock B = (16.3-3.5)/20.55 = 0.62